Find three consecutive odd integers such that the square of the second added to the first is seventy five less than the square of the third.

Answer by checkley77(12844)

Let x, x+2 & x+4 be the 3 integers.
(x+2)^2+x=(x+4)^2-75
x^2+4x+4+x=x^2+8x+16-75
x^2-x^2+5x-8x=-4+16-75
-3x=-63
x=-63/-3
x=21 ans.
Proof:
(21+2)^2+21=(21+4)^2-75
23^2+21=25^2-75
529+21=625-75
550=550